6. Consider an economy that produces and consumes Bread and Automobile. Data for two different years 2005 and 2010 is given in the following table.

 Year 2005 2010

Price of Automobiles $ 5000 $ 6000

Price of a loaf of bread $10 $20

Number of 100 120

automobiles produced

Number of 500,000 400,000

loaves of bread produced

 

 Using the year 2005 as a base year,

a) Calculate the nominal and real GDP of 2010.

b) Find the value of GDP Deflator for the year 2010 and interpret. 

c) Calculate the inflation rate in 2010. 


1
Expert's answer
2021-04-19T07:48:32-0400

a) The nominal GDP of 2010 is:

Yn = 6,000×120 + 20×400,000 = 8,720,000.

The real GDP of 2010 is:

Yn = 5,000×120 + 10×400,000 = 4,600,000.

b) The value of GDP Deflator for the year 2010 is:

Deflator = 8,720,000/4,600,000×100 = 189.56.

So, the real GDP is almost twice lower than the nominal GDP.

c) The inflation rate in 2010 comparing to 2005 was 189.56 - 100 = 89.56%.